Find the equation of the line that passes through points (7,-2) and (5,-1)
Write the equation of the line in slope-intercept form.
I know the the slope intercept for is y = mx +b

Answer by kiru_khandelwal(79)

Equation of the line passing through two points (x1,y1) and (x2,y2) is
(y-y1)/(x-x1) = m
where m = slope of the line = y2-y1/x2-x1
Sp the equation of the line will be
(y-y1)/(x-x1) = (y2-y1)/(x2-x1)...........equation(1)
According to the question x1 = 7 y1 = -2 x2 = 5 and y2 = -1
substituting these values in equation(1)
=> (y-(-2))/(x-7) = (-1-(-2))/(5-7)
=> (y+2)/(x-7) = 1/(-2)
=> 2(y+2) = -(x-7)
=> 2y + 4 = -x + 7
=> 2y = -x + 3
The eqation of line in slope intercept form which is y=mx+c will be
y = -x/2 + 3/2